# VaR at Its Extremes: Impossibilities and Conditions for One-Sided Random Variables

### v2

@@ 4 Further Generalizations and Remarks
The results established in Sections 2 and 3 extend naturally to random variables whose supports begin at arbitrary finite lower end-points
$a_{i}=\sup\{x\in\mathbb{R}:F_{X_{i}}(x)\leq 0\}>-\infty,\qquad\forall i\in\{1,\dots,n\}.$
We denote the corresponding random vector by
$\bm{X}^{\bm{a}}=(X_{1}^{a_{1}},\dots,X_{n}^{a_{n}}),\qquad\bm{a}=(a_{1},\dots,a_{n})\in\mathbb{R}^{n},$
with $\bm{X}^{\bm{0}}=\bm{X}$ representing the previously studied case of random variables supported on $0,\infty)$.
Similarly, the theory extends to random variables possessing arbitrary finite upper end-points
$b_{i}=\inf\{x\in\mathbb{R}:F_{X_{i}}(x)\geq 1\}<\infty,\qquad\forall i\in\{1,\dots,n\},$
for which we write
$\bm{X}^{\bm{b}}=(X_{1}^{b_{1}},\dots,X_{n}^{b_{n}}),\qquad\bm{b}=(b_{1},\dots,b_{n})\in\mathbb{R}^{n}.$
For notational convenience, define the corresponding sum random variables by
$S^{\bm{a}}=\sum_{i=1}^{n}X_{i}^{a_{i}},\qquad S^{\bm{b}}=\sum_{i=1}^{n}X_{i}^{b_{i}}.$
It is immediate that both transformed vectors admit the simple representations
$\bm{X}^{\bm{a}}=\bm{a}+\bm{X},\qquad\bm{X}^{\bm{b}}=\bm{b}-\bm{X},$
and, due to the translation and scale equivariance of VaR, their components satisfy
${\mathrm{VaR}}_{p}[X_{i}^{a_{i}}]$ $={\mathrm{VaR}}_{p}[a_{i}+X_{i}]=a_{i}+{\mathrm{VaR}}_{p}[X_{i}],$
${\mathrm{VaR}}_{p}[X_{i}^{b_{i}}]$ $={\mathrm{VaR}}_{p}[b_{i}-X_{i}]=b_{i}-{\mathrm{VaR}}_{1-p}[X_{i}].$
Both facts will be crucial in what follows.
@@
The preceding proposition highlights an important structural limitation: VaR sub-additivity cannot occur for random variables with finite lower end-points, while VaR super-additivity cannot occur for random variables with finite upper end-points.
Corollary 4.2.
For compactly supported random variables $\bm{X}^{\bm{a},\bm{b}}$, i.e. random variables possessing both finite lower and upper end-points, VaR sub-additivity and VaR super-additivity are each equivalent to VaR additivity. Consequently, such random variables can never exhibit strict VaR sub-additivity or strict VaR super-additivity.
Proposition 4.3.
  • (i)
    and that the function
    is SD. Then $\bm{X}^{\bm{a}}$ is VaR super-additive.
  • (ii)
    and that the function
    is SD. Then $\bm{X}^{\bm{b}}$ is VaR sub-additive.
@@
Using the results we obtained in Proposition 4.3, we can now delineate the sufficient conditions that parallel those of Propositions 3.5 and 3.6. These conditions are easily verifiable and ensure that $\bm{X}^{\bm{a}}$ (resp. $\bm{X}^{\bm{b}}$) is VaR super-additive (resp. VaR sub-additive).
Proposition 4.4.
  • (i)
    is non-increasing, then $\bm{X}^{\bm{a}}$ is VaR super-additive.
  • (ii)
    is non-increasing, then $\bm{X}^{\bm{b}}$ is VaR sub-additive.
@@
    In this setting, the $n$-box $x_{1},b_{1}]\times\dots\times[x_{n},b_{n}]$ contains a “reversed” $n$-simplex with origin $(b_{1},\dots,b_{n})$ and vertices
@@
5 Conclusions
This paper provides a comprehensive characterization of the extremal aggregation behavior of Value-at-Risk for sums of one-sided random variables. We first established an impossibility result: for risks supported on $0,\infty)$, VaR sub-additivity can arise only through exact additivity – a phenomenon exclusive to co-monotonic vectors. On the opposite end of the spectrum, we developed a general and flexible framework for full VaR super-additivity. The key insight is that super-additivity does not follow from dependence or marginal structure in isolation, but from their joint interaction as captured by the NSD and SD conditions. These conditions unify and extend existing results in the literature, while accommodating non-identical margins and a diverse range of negative dependence structures.
We further showed that the theory remains robust under shifts, reflections, and monotone convex transformations of the components, and that analogous principles govern aggregation when the random variables have arbitrary finite endpoints. Taken together, the results reveal a sharp dichotomy: in lower-bounded settings, VaR is structurally incompatible with sub-additivity yet naturally exhibits super-additivity under suitable dependence–margin configurations, whereas the pattern is reversed in upper-bounded settings. This characterization not only clarifies the conditions under which VaR behaves as a diversification-averse or diversification-seeking risk measure, but also offers practical criteria for detecting such behavior in applications involving heavy tails or negatively dependent risks.

### v3

@@ 4 Further Generalizations and Remarks
In many actuarial applications, losses are sometimes bounded below by a positive constant $a_{i}$. A canonical example is the franchise deductible (Klugman et al., , 2019), under which the full loss is paid only when it exceeds a fixed threshold, effectively shifting the support to $a_{i},\infty)$. The results of Sections 2 and 3 extend naturally to random variables with arbitrary finite lower endpoints
$a_{i}=\sup\{x\in\mathbb{R}:F_{X_{i}}(x)\leq 0\}>-\infty,\qquad\forall i\in\{1,\dots,n\},$
and, symmetrically, to random variables with arbitrary finite upper endpoints
$b_{i}=\inf\{x\in\mathbb{R}:F_{X_{i}}(x)\geq 1\}<\infty,\qquad\forall i\in\{1,\dots,n\}.$
We denote the corresponding shifted and reflected random vectors and their aggregate sums by
$\bm{X}^{\bm{a}}=(X_{1}^{a_{1}},\dots,X_{n}^{a_{n}}),\quad\bm{a}=(a_{1},\dots,a_{n})\in\mathbb{R}^{n},\qquad\bm{X}^{\bm{b}}=(X_{1}^{b_{1}},\dots,X_{n}^{b_{n}}),\quad\bm{b}=(b_{1},\dots,b_{n})\in\mathbb{R}^{n},$
$S^{\bm{a}}=\sum_{i=1}^{n}X_{i}^{a_{i}},\qquad S^{\bm{b}}=\sum_{i=1}^{n}X_{i}^{b_{i}},$
where $\bm{X}^{\mathbf{0}}=\bm{X}$ recovers the previously studied case of risks supported on $0,\infty)$. As the following remark makes clear, all results in this section are consequences of the translation and reflection equivariance of VaR; they are stated explicitly in the shifted and reflected parameterizations for the reader’s convenience and for direct applicability in actuarial practice.
Remark 4.1.
Since $X_{i}^{a_{i}}=a_{i}+X_{i}$ with $X_{i}\geq 0$, the shifted vector $\bm{X}^{\bm{a}}$ reduces to the zero-endpoint case via ${\mathrm{VaR}}_{p}[X_{i}^{a_{i}}]=a_{i}+{\mathrm{VaR}}_{p}[X_{i}]$. Similarly, $X_{i}^{b_{i}}=b_{i}-X_{i}$ and the reflection identity ${\mathrm{VaR}}_{p}[b_{i}-X_{i}]=b_{i}-{\mathrm{VaR}}_{1-p}[X_{i}]$ reduce VaR super-additivity (resp. VaR sub-additivity) of $\bm{X}^{\bm{b}}$ to VaR sub-additivity (resp. VaR super-additivity) of $\bm{X}$, handled by Section 2 (resp. Section 3). The propositions below make these reductions explicit in parameterizations that arise naturally in practice.
@@
The proof is relegated to Appendix Ai.
The preceding proposition highlights an important structural contrast: VaR sub-additivity cannot occur for random variables with finite lower endpoints, while VaR super-additivity cannot occur for random variables with finite upper endpoints.
Remark 4.3.
For compactly supported random variables $\bm{X}^{\bm{a},\bm{b}}$ i.e. those with finite lower and upper endpoints, Proposition 4.2 implies that VaR sub-additivity and VaR super-additivity are each equivalent to VaR additivity. Consequently, strict forms of either property are impossible in this setting. This conclusion aligns with the result established for integrable random variables in Imamura and Kato, (2025).
Proposition 4.4.
  • (i)
    and that the function
    is SD. Then $\bm{X}^{\bm{a}}$ is VaR super-additive.
  • (ii)
    and that the function
    is SD. Then $\bm{X}^{\bm{b}}$ is VaR sub-additive.
The complete proof appears in Appendix Aii.
Using the results we obtained in Proposition 4.4, we can now delineate the sufficient conditions that parallel those of Propositions 3.6 and 3.7. These conditions are easily verifiable and ensure that $\bm{X}^{\bm{a}}$ (resp. $\bm{X}^{\bm{b}}$) is VaR super-additive (resp. VaR sub-additive).
Proposition 4.5.
  • (i)
    is non-increasing, then $\bm{X}^{\bm{a}}$ is VaR super-additive.
  • (ii)
    is non-increasing, then $\bm{X}^{\bm{b}}$ is VaR sub-additive.
The result is proved in Appendix Aiii.
Proposition 4.6.
  • (i)
  • (ii)
The proof is provided in Appendix Aiv
5 Conclusions
This paper provides a comprehensive characterization of the extremal aggregation behavior of Value-at-Risk for sums of one-sided random variables. We first established an impossibility result that is a direct extension of the finding in Imamura and Kato, (2025). For risks supported on $0,\infty)$, possibly non-integrable, VaR sub-additivity can arise only through exact additivity – a phenomenon exclusive to co-monotonic vectors. On the opposite end of the spectrum, we developed a general and flexible framework for full VaR super-additivity. The key insight is that super-additivity does not follow from dependence or marginal structure in isolation, but from their joint interaction as captured by the NSD and SD conditions. These conditions unify and extend existing results in Chen and Shneer, (2026), while accommodating non-identical margins and a diverse range of negative dependence structures.
We further showed that the theory remains robust under monotone convex transformations of the components, and that analogous principles govern aggregation when the random variables have arbitrary finite endpoints. Taken together, the results reveal a sharp dichotomy: in lower-bounded settings, VaR is structurally incompatible with sub-additivity yet naturally exhibits super-additivity under suitable dependence–margin configurations, whereas the pattern is reversed in upper-bounded settings. This characterization not only clarifies the conditions under which VaR behaves as a diversification-averse or diversification-seeking risk measure, but also offers practical criteria for detecting such behavior in applications involving heavy tails or negatively dependent risks.
Appendix A Proofs of Section 4
i Proof of Proposition 4.2
@@ Appendix A Proofs of Section 4
ii Proof of Proposition 4.4
@@
iii Proof of Proposition 4.5
@@
    In this setting, the $n$-box $x_{1},b_{1}]\times\dots\times[x_{n},b_{n}]$ contains the “reversed” $n$-simplex with origin $(b_{1},\dots,b_{n})$ and vertices